OpenStudy (gabbyalicorn):

Which ordered pair is the solution to the system of equations? x+y=-1 x-3y=11 A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (lochana):

write equations for y.

OpenStudy (lochana):

you can re-arrange equations. \[y = -1 -x \]\[y = \frac{x-11}{3}\]

OpenStudy (gabbyalicorn):

but -1-x =−x−1 & x-11/3 is justx+−113

OpenStudy (gabbyalicorn):

-11/3 i don't get it

OpenStudy (lochana):

no. we are asked to find solutions for given 2 equations. that means you need to find the point where those equations meet

OpenStudy (lochana):

okay?

OpenStudy (gabbyalicorn):

Oh, that makes more sense

OpenStudy (lochana):

yes. so at the common point x and y of each equation are same that means y points are equal

OpenStudy (lochana):

so we equate \[-1 -x = \frac{x-11}{3}\]

OpenStudy (lochana):

\[3(-1-x) = x - 11\]\[4x = 8\]x = 2

OpenStudy (gabbyalicorn):

that would be 4/3 for the first one

OpenStudy (lochana):

yes. you are correct

OpenStudy (gabbyalicorn):

second one is 2

OpenStudy (gabbyalicorn):

:D

OpenStudy (lochana):

bye

OpenStudy (gabbyalicorn):

wait!!

OpenStudy (gabbyalicorn):

so its 2,-3?